Solve the following equations. 
t/5 = 10
Solve the following equations. 
t/5 = 10

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

t/5=10
t = 10 × 5
t=50

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve this question as follows :
T/5 = 10
T = 5 × 10
T = 50
 
Hope it helps !
Thank you ✌️